1 cup = 16 tablespoons
1 tablespoon = 3 teaspoons

A certain recipe calls for ¼ cup of white sugar, ½ cup of brown sugar, 2 ¼ cups of flour, ¼ teaspoon of salt, and 2 tablespoons of cornstarch. The amount of cornstarch called for is approximately what fraction of the total amount of sugar called for?

Answer by ptaylor(1894)


Lets deal in tablespoons
white sugar=1/4 cup=(1/4)*16=4 tablespoons
brown sugar=1/2 cup=(1/2)*16=8 tablespoons
Total amount of sugar called for=4+8=12 tablespoons
Amount of cornstarch called for=2 tablespoons
cornstarch = 2/12=1/6 of the total amount of sugar called for
Hope this helps---ptaylor